What are the magnitude and direction of the electric field whose force will balance the weight of the following? (Use the information in the table below).

Particle Charge (C) Mass (kg)
Electron (e) −1.6021765 ✕ 10−19 9.1094 ✕ 10−31
Proton (p) +1.6021765 ✕ 10−19 1.67262 ✕ 10−27
Neutron (n) 0 1.67493 ✕ 10−27

(a)

an electron

(b)

a proton
Answer #1

for electron

qE = mg

1.6* 10^-19 E = 9.11* 10^-31* 9.8

E = 5.579* 10^-11 N/C

directed upwards

=====

for proton

qE = mg

E = 1.67* 10^-27* 9.8 / ( 1.6* 10^-19)

E = 1.0229* 10^-7 N/C

directed downwards